Given a required return of 10%p.a., consider the following investments:

                                     Cash Flow

Year                             project1                     project2

0                                -$40000                         -$6300

1                                 $20000                          $3500

2                                 $20000                          $3500

3                                  $20000                          $3500

a) Calculate the profitability index (PI) of the above projects.
b) Calculate the NPV of the above projects.
c) Which of these projects out to be selected if the firm can only select one of them?

Please calculate the process
Answer #1

(a) Project 1:

Initial Cash Flow = - $ 40000 and Annual Cash Flows = $ 20000, Tenure = 3years, Required Return = 10 %

Present Value of Cash Flows = PV1 = 20000 x (1/0.1) x [1-{1/(1.1)^(3)}] = $ 49737.0398

Profitability Index (PI) = 49737.0398 / 40000 ~ 1.24

Project 2:

Initial Cash Flow = - $ 6300 and Annual Cash Flows = $ 3500, Tenure = 3years, Required Return = 10 %

Present Value of Cash Flows = PV1 = 3500 x (1/0.1) x [1-{1/(1.1)^(3)}] = $ 8703.982

Profitability Index (PI) = 8703.982 / 6300 ~ 1.38

(b) Project 1 NPV = 49737.0398 - 40000 = $ 9737.0398

Project 2 NPV = 8703.982 - 6300 = $ 2403.9819 ~ $ 2403.98

(c) The project to be selected should be Project 1, as the same provides a greater positive NPV (greater positive value addition) to the entity undertaking the project, even though project 1's PI is lower than that of Project 2.